OpenStudy (anonymous):

expand the binomial = (x+1)^5

OpenStudy (sidsiddhartha):

use (x + y)5 = x^5 +5x^4 y + 10x^3 y^2 +10x^2 y^3 +5xy^4 + y^5 put y=1

OpenStudy (anonymous):

is this a identity?

OpenStudy (anonymous):

how did you do that?

OpenStudy (sidsiddhartha):

it's a binomial expantion http://en.wikipedia.org/wiki/Binomial_theorem

OpenStudy (anonymous):

by using pascal's triangle.

OpenStudy (anonymous):

http://www.mathsisfun.com/pascals-triangle.html

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome .